# Topological orbital Hall effect caused by skyrmions and antiferromagnetic skyrmions

### v2

@@ Abstract
Abstract. The topological Hall effect is a hallmark of topologically non-trivial magnetic textures such as magnetic skyrmions. It quantifies the transverse electric current once an electric field is applied and occurs as a consequence of the emergent magnetic field of the skyrmion. Likewise, an orbital magnetization is generated. Here we show that the charge currents are orbital polarized even though the conduction electrons couple to the skyrmion texture via their spin. The topological Hall effect is accompanied by a topological orbital Hall effect even for $s$ electrons without spin-orbit coupling. As we show, antiferromagnetic skyrmions and antiferromagnetic bimerons that have a compensated emergent field, exhibit a topological orbital Hall conductivity that is not accompanied by charge transport and can be orders of magnitude larger than the topological spin Hall conductivity. Skyrmionic textures serve as generators of orbital currents that can transport information and give rise to considerable orbital torques.
@@ Introduction
Over the recent years, the orbital degree of freedom has become increasingly relevant in the field of spinorbitronics, manifesting itself in phenomena such as orbital magnetization [40, 41, 42, 43, 44, 45], orbital torque [46, 47], orbital Edelstein effect [48, 49] and orbital Hall effect [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]. In the presence of a magnetic field, charge currents form circular trajectories and generate an orbital angular momentum that can be calculated via the modern formulation of orbital magnetization [40, 41, 42, 43, 44, 45]. This orbital magnetization has been considered for quantum Hall systems as well as skyrmions [78, 79, 45]. Recently, we have shown that the skipping orbits in quantum Hall systems lead to orbital polarized edge currents [67] that give rise to an orbital Hall effect accompanying the (charge) Hall effect once a transverse electric field is applied.
@@
In this paper, we show that skyrmions give rise to a topological orbital Hall effect (Fig. 1). The orbital currents appear additionally to the topological (charge) Hall effect giving rise to transverse orbital polarized currents. By using a tight-binding model and a Berry curvature approach, we systematically compare the charge, spin and orbital Hall conductivities. Furthermore, we analyze the edge states that occur when the skyrmion is on the nanometer scale. In this scenario the emergent field of the skyrmion is extremely large and Landau levels form. As we show, when two oppositely magnetized skyrmions are coupled to form an antiferromagnetic skyrmion, the topological Hall effect is compensated and only orbital and spin Hall effects emerge. Since the orbital angular momentum of the conduction electrons in a skyrmion system can be arbitrarily high, the topological orbital Hall conductivity can be orders of magnitude larger than the topological spin Hall conductivity.
@@
While the two blocks in the band structure are characterized by opposite spin alignment, that is rather homogeneous within a block ($s_{\nu,z}(\mathbf{k})\approx\pm\hbar/2\,\braket{m_{z}}$ with $\braket{m_{z}}$ the average out-of-plane magnetic moment of the skyrmion), the orbital angular momentum $L_{\nu,z}(\mathbf{k})$ is also roughly opposite comparing the two blocks but changes within a block. We have added it as a color code to Fig. 2(a). Starting from the lowest band, it is negative and increases in magnitude until it changes sign near $E=-m=-5\,\mathrm{eV}$ and decreases back to zero. Its magnitude is much larger than the spin by multiple $\hbar$ (cf. legend in Fig. 2a).
@@
First, and most importantly, the orbital Hall conductivity (panel c) is non-zero. This means, a skyrmion exhibits orbital and spin polarized currents; a topological orbital Hall effect is superimposed on top of the previously observed topological Hall and topological spin Hall effects.
@@
Within a block, the topological Hall conductivity (b) and spin Hall conductivity change sign (d). This is because the electrons in the lower half of a block are characterized by a positive effective mass and in the upper half by a negative effective mass (hole-like behavior) which is determined by the band structure of the underlying square lattice. Particles of opposite mass are deflected into opposite directions by the same emergent magnetic field. However, by changing the sign of the effective mass, $L_{z}$ changes sign as well, as in classical physics. For this reason, the orbital Hall conductivity (c) remains always positive.
@@ Methods
Here, $\mathbf{v}=\frac{1}{\hbar}\nabla_{\mathbf{k}}H$ is the velocity operator. $L_{\nu,z}(\mathbf{k})$ are the diagonal elements of the tensor $L_{\nu,z}(\mathbf{k})=\langle\nu\mathbf{k}|L_{z}|\nu\mathbf{k}\rangle$. Note that we have corrected a mistake in Ref. [57] where $\mathrm{Im}$ was used instead of the imaginary unit $i$.

### v3

@@ Abstract
Abstract. The topological Hall effect is a hallmark of topologically non-trivial magnetic textures such as magnetic skyrmions. It quantifies the transverse electric current that is generated once an electric field is applied and occurs as a consequence of the emergent magnetic field of the skyrmion. Likewise, an orbital magnetization is generated. Here we show that the charge currents are orbital polarized even though the conduction electrons couple to the skyrmion texture via their spin. The topological Hall effect is accompanied by a topological orbital Hall effect even for $s$ electrons without spin-orbit coupling. As we show, antiferromagnetic skyrmions and antiferromagnetic bimerons that have a compensated emergent field, exhibit a topological orbital Hall conductivity that is not accompanied by charge transport and can be orders of magnitude larger than the topological spin Hall conductivity. Skyrmionic textures serve as generators of orbital currents that can transport information and give rise to considerable orbital torques.
@@ Introduction
Over the recent years, the orbital degree of freedom has become increasingly relevant in the field of spinorbitronics, manifesting itself in phenomena such as orbital magnetization [40, 41, 42, 43, 44, 45], orbital torque [46, 47], orbital Edelstein effect [48, 49] and orbital Hall effect [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77]. In the presence of a magnetic field, charge currents form circular trajectories and generate an orbital angular momentum that can be calculated via the modern formulation of orbital magnetization [40, 41, 42, 43, 44, 45]. This orbital magnetization has been considered for quantum Hall systems as well as skyrmions [78, 79, 45]. Recently, we have shown that the skipping orbits in quantum Hall systems lead to orbital-polarized edge currents [67] that give rise to an orbital Hall effect accompanying the (charge) Hall effect once a transverse electric field is applied.
@@
In this paper, we show that skyrmions give rise to a topological orbital Hall effect (Fig. 1). The orbital currents appear additionally to the topological (charge) Hall effect giving rise to transverse orbital-polarized currents. By using a tight-binding model and a Berry curvature approach, we systematically compare the charge, spin and orbital Hall conductivities. Furthermore, we analyze the edge states that occur when the skyrmion is on the nanometer scale. In this scenario the emergent field of the skyrmion is extremely large and Landau levels form. As we show, when two oppositely magnetized skyrmions are coupled to form an antiferromagnetic skyrmion, the topological Hall effect is compensated and only orbital and spin Hall effects emerge. Since the orbital angular momentum of the conduction electrons in a skyrmion system can be arbitrarily high, the topological orbital Hall conductivity can be orders of magnitude larger than the topological spin Hall conductivity.
@@
While the two blocks in the band structure are characterized by opposite spin alignment, that is rather homogeneous within a block ($S_{\nu,z}(\mathbf{k})\approx\pm\hbar/2\,\braket{m_{z}}$ with $\braket{m_{z}}$ the average normalized out-of-plane magnetic moment of the skyrmion), the orbital angular momentum $L_{\nu,z}(\mathbf{k})$ is also roughly opposite comparing the two blocks but changes within a block. We have added it as a color code to Fig. 2(a). Starting from the lowest band, it is negative and increases in magnitude until it changes sign near $E=-m=-5\,\mathrm{eV}$ and decreases back to zero. Its magnitude is much larger than the spin by multiple $\hbar$ (cf. legend in Fig. 2a).
@@
First, and most importantly, the orbital Hall conductivity (panel c) is non-zero. This means, a skyrmion exhibits orbital- and spin-polarized currents; a topological orbital Hall effect is superimposed on top of the previously observed topological Hall and topological spin Hall effects.
@@
Within a block, the topological Hall conductivity (b) and spin Hall conductivity (d) change sign. This is because the electrons in the lower half of a block are characterized by a positive effective mass and in the upper half by a negative effective mass (hole-like behavior) which is determined by the band structure of the underlying square lattice. Particles of opposite mass are deflected into opposite directions by the same emergent magnetic field. However, by changing the sign of the effective mass, $L_{z}$ changes sign as well, as in classical physics. For this reason, the orbital Hall conductivity (c) remains always positive.
@@ Methods
Here, $\mathbf{v}=\frac{1}{\hbar}\nabla_{\mathbf{k}}H$ is the velocity operator. $L_{\nu,z}(\mathbf{k})$ are the diagonal elements of the tensor $L_{\nu,z}(\mathbf{k})=\langle\nu\mathbf{k}|L_{z}|\nu\mathbf{k}\rangle$. Note that we have corrected a mistake in Ref. [57] where $\mathrm{Im}$ was used instead of the imaginary unit $\mathrm{i}$.